### v1

# Error Bounds of Supervised Classification from Information-Theoretic Perspective

Abstract
There remains a list of unanswered research questions on deep learning (DL), including the remarkable generalization power of overparametrized neural networks, the efficient optimization performance despite the non-convexity, and the mechanisms behind flat minima in generalization. In this paper, we adopt an information-theoretic perspective to explore the theoretical foundations of supervised classification using deep neural networks (DNNs). Our analysis introduces the concepts of fitting error and model risk, which, together with generalization error, constitute an upper bound on the expected risk. We demonstrate that the generalization errors are bounded by the complexity, influenced by both the smoothness of distribution and the sample size. Consequently, task complexity serves as a reliable indicator of the dataset’s quality, guiding the setting of regularization hyperparameters. Furthermore, the derived upper bound fitting error links the back-propagated gradient, Neural Tangent Kernel (NTK), and the model’s parameter count with the fitting error. Utilizing the triangle inequality, we establish an upper bound on the expected risk. This bound offers valuable insights into the effects of overparameterization, non-convex optimization, and the flat minima in DNNs.Finally, empirical verification confirms a significant positive correlation between the derived theoretical bounds and the practical expected risk, confirming the practical relevance of the theoretical findings.
keywords:
deep learning, generalization, Kullback-Leibler (KL) divergence, information theory.
1 Introduction
Deep learning (DL)’s remarkable practical achievements have a profound impact on the conceptual foundations of machine learning and artificial intelligence. Unraveling the factors that contribute to the performance of deep neural networks (DNNs) would not only enhance model interpretability but also facilitate the development of more principled and trustworthy architectural designs. Despite the impressive empirical success, explaining why DNNs can generalize well despite their overwhelming capacity remains an open research challenge. In the statistical learning framework, the objective of learning is to minimize the expected risk, which is composed of the empirical risk and the generalization error. A model is considered effective if both generalization error and empirical risk are minimized. The empirical risk refers to the error or loss of a model when evaluated on the training dataset. The generalization error measures the population risk, representing the average loss over new, randomly sampled data from the population. Extensive experimental evidence indicates that generalization of DNNs significantly influences prediction performance [1], alongside their expressivity [2, 3] and trainability [4, 5]. Some classical theory work attributes the generalization ability to the employment of a low-complexity class of hypotheses and has proposed various complexity measures, including the Vapnik-Chervonenkis (VC) dimension[6], Rademacher complexity[7], and uniform stability[8, 9], to control generalization error. However, recent works like Zhang et al. [10], Zhang et al. [11], He et al. [12] and Belkin et al. [13] revel that high-capacity DNNs can still generalize effectively with genuine labels, yet generalize poorly when trained with random labels. This finding contradicts the traditional wisdom in statistical learning theory, which posits that complex models tend to overfit the training data[14, 6, 7].
In addition to the complexity of the hypothesis space, research[15, 16] highlights the impact of optimization methods, training data, and neural network architectures on the generalization of deep models.
In recent years, information-theoretic approaches, which jointly consider the hypothesis class, learning algorithm, and the data distribution, have drawn considerable attention since the works of Russo and Zou [17] and Xu and Raginsky [18]. The advantage of applying information-theoretic tools in understanding generalization lies in the fact that generalization is influenced by various aspects, including the data distribution, the complexity of the hypothesis class, and the characteristics of the optimization algorithm. For instance, the mutual information (MI) between the training sample and the trained parameter weights has been applied in various generalization bounds [19, 20, 21, 22].
(a) Classical bound and our proposed bound on expected risk.
(b) The logical framework underpinning our bound.
Figure 1: The function $U(\cdot)$ depicted in the Figure1b denotes an upper bound. Our proposed upper bound decomposes the expected error into three distinct components: model risk, fitting error, and generalization error, by utilizing the triangle inequality. Subsequently, we perform a separate analysis of the factors influencing the upper bounds of both the fitting error and the generalization error.
Existing error bounds, grounded in information theory, primarily concentrate on bounding generalization errors, with limited examination of empirical and expected risks. To the best of our knowledge, these bounds fall short in explaining over-parameterization and non-convex optimization mechanisms in DNNs. In addition, mutual information-based upper bounds encounter significant computational challenges and often impractical for real-world application.
The ability of DNNs to achieve high performance despite their immense capacity remains an unresolved question in the field. Consequently, this paper investigates the error bounds of supervised classification, a central topic in the realm of machine learning. Classical statistical learning theory decomposes expected risk into empirical risk and generalization error, primarily concentrating on bounding the latter. In contrast, our approach extends this framework by introducing the concept of fitting error alongside model risk. These components together with the generalization error form an upper bound on the expected risk, offering a more comprehensive perspective than traditional focus on generalization error alone. We demonstrate that generalization errors can be effectively controlled by the complexity of the task, which is influenced by both the support size of the data distribution and the sample size. With a sufficiently large sample size, we establish a correlation between the complexity and the smoothness of the data distribution, suggesting that the complexity of the task can serve as an indicator of the training dataset reliability and guide the setting of regularization hyperparameters. In addition, we derive an upper bound on the fitting error, revealing a connection between the back-propagated gradient, NTK, and the model parameter count to the fitting error. This analysis sheds light on the role of over-parameterization and non-convex optimization in DNNs. In our proofs, we remove the dependence of the error bounds on the specific expression of the loss function based on the Kullback-Leibler (KL) divergence between the empirical probability mass function (PMF) and the real probability mass function (PMF), which is also central to our methodology. Finally, by integrating complexity and fitting error, we utilize the triangle inequality to establish an upper bound on the expected risk. This bound highlights the impact of complexity of the learning task and the fitting errors on the expected risk, and under mild conditions, it yields several important insights:
  • Increasing training samples and decreasing the maximum the loss value contribute to reduction in the expected risk.
  • The expected risk can be effectively controlled by the gradient of the Kullback-Leibler (KL) divergence between the empirical conditional probability distribution and the predictive probability distribution of the model.
  • Gradient-based algorithms can lead to lower expected risk as the number of parameters grows.
  • A smaller maximum eigenvalue of equal-input Neural Tangent Kernel (eNTK), $\lambda_{\max}(H(f_{\theta}(x)))$, which is proved to be equivalent to the flat minima, is advantageous in minimizing the expected risk.
The article’s structure is organized as follows. In Section 3 we begin by defining the fundamental concepts and objects employed throughout the paper to establish a clear foundation. In Section 2, we delves into the existing literature, discuss relevant research. In Section 4 we introduce the concept of complexity and prove that generalization errors can be bounded by measure, emphasizing its significance in learning. In Section 5, an upper bound on fitting error is derived, providing further insights in the optimization process and the role of model parameters. In Section 6, we derive the an upper bound on expected error and analyse the learning techniques for DNNs based on this bound. In Section 7, empirical evidence is presented, demonstrating a strong positive correlation between the derived upper bounds and the expected risk in real-world scenarios, thereby validating our theoretical findings. In Section 8, the article concludes with a summary of our key contributions, and highlights some important implications.
2 Related Work
Our work is intrinsically connected to the extensive literature on information-theoretic generalization bounds, which have been empirically demonstrated their efficacy in capturing the generalization behavior of Deep Neural Networks (DNNs) [23, 24]. The original information-theoretic bound introduced by [18] has been extended or improved in various ways, such as the chaining method [21, 25, 26] , the random subset and individual technique [23, 27, 28, 29, 30].
Information-theoretic generalization bounds have also been explored for some specific algorithms. For example, the-weight based information-theoretic bounds have been successfully applied to characterize the generalization properties of the stochastic gradient Langevin dynamics (SGLD) algorithm [20, 27, 23, 28, 29, 31], and more recently, these bounds have also been used to analyze either vanilla Stochastic Gradient Descent (SGD) [32, 33] or the stochastic differential equations (SDEs) approximation of SGD [33].
The information bottleneck principle [34, 35, 36] , from the perspective of the information change in neural networks, posits for better prediction, neural network learns useful information and filters out redundant information.
Beyond supervised learning, information-theoretic bounds are also useful in a variety of learning scenarios, including meta-learning [37], semi-supervised learning[38], transfer learning[39], among others.
Another popular claim in the literature is that flat minima generalize better than sharp (non-flat) minima [40, 15]. In response to this notion, Foret et al. [41] proposed Sharpness-Aware Minimization (SAM), an approach designed to directly optimize for flat solutions. Flatness is often quantified mathematically with the largest eigenvalue of the Hessian matrix of the loss with respect to the model parameters on the training set [42, 43, 44].
A parallel related line of research is the over-parameterization theory of DNNs [45, 46, 47, 48, 49, 50]. Despite the inherent non-convexity of the objective functions, empirical evidence indicates that gradient-based methods, such as stochastic gradient descent (SGD), are capable of converging to the global minima in these networks. Key insights from Du et al. [45], Chizat et al. [51], Arjevani and Field [52] emphasize the pivotal role of the over-parameterization [53] in enhancing the generalization ability of DNNs. The neural tangent kernel (NTK) thus has emerged as a pivotal concept, as it captures the dynamics of over-parameterized neural network trained by gradient descent [54]. Notably, recent studies [45, 46, 47] demonstrate that, with specialized scaling and random initialization, the dynamics of gradient descent (GD) in continuous-width multi-layer neural networks can be tracked via NTK and the convergence to the global optimum with a linear rate can be proven.
3 Preliminaries
In this section, we provide essential background information on learning algorithms and fundamental concepts, particularly focusing on generalization error. Our analysis on supervised classification is conducted within the standard statistical learning framework. The instance space, denoted as $\mathcal{Z}:=\mathcal{X}\times\mathcal{Y}$, consists of input feature space $\mathcal{X}$ and label set $\mathcal{Y}$. The hypothesis space, a set of functions parameterized by $\Theta$, is represented by $\mathcal{F}_{\Theta}=\{f_{\theta}:\theta\in\Theta\}$, where $\Theta$ denotes the parameter space. Training data is represented by an $n$-tuple $S^{n}=\{(x^{(i)},y^{(i)})\}_{i=1}^{n}$, which consists of independent and identically distributed (i.i.d.) samples drawn from an unknown true distribution $\bar{q}$. The empirical probability mass function (PMF) of $(X,Y)$ derived from these samples $S^{n}$, denoted by $q^{n}$ (or simply $q$), is defined as $q(x,y):=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{x,y\}}(x^{(i)},y^{(i)})$. According to the (weak) law of large numbers, $q$ converges to $\bar{q}$ in probability. We simplify the notation by defining $q_{Y|x}(y):=q_{Y|X}(y|x),q_{Y|x}=(q_{Y|x}(y_{1})$, and $\cdots,q_{Y|x}(y_{|\mathcal{Y}|}))^{\top}$, where $q_{X}(x)$, and $q_{Y|X}(y|x)$ represent the marginal and conditional PMF, respectively, with $y_{i}\in\mathcal{Y}$ for each $i$. We introduce a function $R_{\ell}(f_{\theta},q):\mathcal{F}_{\Theta}(x)\times\Delta^{|\mathcal{Y}|}\to\mathbb{R}$, as expressed as follows:
$R_{\ell}(f_{\theta},q)=\mathbb{E}_{(X,Y)\sim q}\ell(f_{\theta}(x),y).$ (1)
Let $\mathcal{R}(f_{\theta})$ denote the expected risk of a hypothesis $f_{\theta}$ (abbreviation: $f$) under a given loss function $\ell:\mathcal{F}_{\Theta}(x)\times\mathcal{Y}\to\mathbb{R}$, as expressed as follows:
$\mathcal{R}(f_{\theta})$ $=\mathbb{E}_{(X,Y)\sim\bar{q}}\ell(f_{\theta}(x),y)$ (2)
$=R_{\ell}(f_{\theta},\bar{q}).$
The empirical risk, $\mathcal{R}_{S}(f_{\theta})$, is defined as the average loss over the training data:
$\mathcal{R}_{S}(f_{\theta})$ $=\frac{1}{n}\sum_{i=1}^{n}\ell(f_{\theta}(x^{(i)}),y^{(i)})$ (3)
$=\mathbb{E}_{(X,Y)\sim q}\ell(f_{\theta}(x),y)$
$=R_{\ell}(f_{\theta},q).$
For a hypothesis $f_{\theta}\in\mathcal{F}$, the $j$-th element of $f_{\theta}(x)$ is denoted by $[f_{\theta}(x)]_{j}$. To convert this into a probability distribution, we apply the softmax function, resulting in the prediction probabilities for $x$ as follows:
$p_{Y|x}=(\frac{e^{f_{\theta}(x)_{1}}}{Z_{\theta}(x)},\cdots,\frac{e^{f_{\theta}(x)_{|\mathcal{Y}|}}}{Z_{\theta}(x)})^{\top}$ (4)
where $Z_{\theta}(x)$, known as the normalizing constant (a.k.a. the partition function), is given by:
$Z_{\theta}(x)=\sum_{j=1}^{|\mathcal{Y}|}e^{f_{\theta}(x)_{j}}.$ (5)
The model risk, $\mathcal{R}_{M}(f_{\theta})$, is defined as
$\mathcal{R}_{M}(f_{\theta})$ $=\mathbb{E}_{(X,Y)\sim p}\ell(f_{\theta}(x),y)$ (6)
$=R_{\ell}(f_{\theta},p).$
where $p(x,y)=q_{X}(x)p_{Y|x}(y)$. For a cross-entropy loss. $\mathcal{R}_{M}(f_{\theta})$ simplifies to negative conditional entropy of random variables $X$ and $Y$ when $(X,Y)\sim p$:
$\mathcal{R}_{M}(f_{\theta})=-\mathbb{E}_{(X,Y)\sim p}\log p_{Y|x}=H_{p}(Y|X),$ (7)
where $H_{p}(Y|X)$ represents the conditional entropy of $Y$ given $X$.
Sanov’s theorem [55] states that:
$I_{p_{Y|x}}(q_{Y|x})$ $=\lim_{n\to\infty}-\frac{1}{n}\log\Pr(q_{Y|x}|p_{Y|x})$ (8)
$=D_{KL}(q_{Y|x}\|p_{Y|x})$
, where $I_{p_{Y|x}}(q_{Y|x})$ is called the rate function, representing the probability that the empirical PMF of $p_{Y|x}$ is $q_{Y|x}$. Motivated by this, we introduce the expected rate function (ERF), defined as
$\mathcal{R}_{ERF}(\theta,q)$ $:=\mathbb{E}_{X\sim q_{X}}D_{KL}(q_{Y|x}\|p_{Y|x}).$ (9)
Substituting the parameterized form $p_{Y|x}$ into $\mathcal{R}_{ERF}(\theta,q)$ yields:
$\mathcal{R}_{ERF}(\theta,q)$ $=\mathbb{E}_{X\sim q_{X}}[\log Z_{\theta}(x)-q_{Y|x}^{\top}f_{\theta}(x)]-H_{q}(Y|X)$ (10)
$=R_{SCE}(f_{\theta},q)-H_{q}(Y|X),$
where $(X,Y)\sim q$, SCE stands for softmax cross entropy. Since $H_{q}(Y|X)$ is independent of the model $f_{\theta}$, then during optimization process, $\mathcal{R}_{ERF}(\theta,q)$ is equivalent to $R_{SCE}(f_{\theta},q)$. To clarify the relationships between expected risk, empirical risk, model risk, and generalization error, we will utilize the function $R_{\ell}(f_{\theta},\cdot)$ to represent these risks and errors in the subsequent analysis. In classification tasks, the objective is typically to find a hypothesis $f_{\theta}$ to minimize the incompatible expected risk $\mathcal{R}(f_{\theta})$. This often achieved by minimizing the computable empirical risk $\mathcal{R}_{S}(f_{\theta})$ (i.e., empirical risk minimization). One goal of generalization theory is to explain and justify when and how minimizing $\mathcal{R}_{S}(f_{\theta})$ is a sensible approach to minimizing $\mathcal{R}(f_{\theta})$ by analyzing the generalization error:
$\mathrm{gen}(\ell,f_{\theta})$ $:=|R_{\ell}(f_{\theta},\bar{q})-R_{\ell}(f_{\theta},q)|.$ (11)
Similar to the generalization error, we introduce a new concept, the fitting error, to measure the gap between empirical risk and model risk
$\mathrm{fit}(\ell,f_{\theta}):=|R_{\ell}(f_{\theta},p)-R_{\ell}(f_{\theta},q)|.$ (12)
The fitting error essentially reflects the discrepancy between the distribution produced by the model and the empirical PMF. To focus on the effect of the characteristics of the model, we define the normalized fitting error, denoted by :
${\mathrm{fit_{n}}(\ell,f_{\theta})}:={\mathrm{fit}(\ell,f_{\theta})}/{\sqrt{\mathbb{E}_{X}[\|\ell(f_{\theta}(x))\|^{2}_{2}]}},$ (13)
where $\ell(f_{\theta}(x))=(\ell(f_{\theta}(x),y_{1}),\cdots,\ell(f_{\theta}(x),y_{|\mathcal{Y}|}))^{\top}$. We also introduce the equal-input neural tangent kernel (eNTK), which is the special case of empirical neural tangent kernel (NTK), which operates between input feature $x$ and itself[54]:
$H(f_{\theta}(x))=\nabla_{\theta}f_{\theta}(x)^{\top}\nabla_{\theta}f_{\theta}(x).$ (14)
The theorems proved in the subsequent sections will rely on the following lemma.
Lemma 1 (Markov inequality).
Let $Z$ be a non-negative random variable. Then, for all $t\geq 0$, $\operatorname{Pr}(Z\geq t)\leq\mathbb{E}[Z]/t$.
4 Generalization and Complexity of task from the lens of information theory
This section aims to establish the connection between generalization error and the complexity of the classification tasks, and explore the application of complexity in practical classification scenarios. Some classical theoretical studies attribute the generalization ability to the employment of a low-complexity hypotheses and have introduced various complexity measures for hypotheses to control generalization error. The source of $\mathrm{gen}(\ell,f_{\theta})$ is caused by the insufficiency of training samples, which fail to adequately represent the distribution of the entire dataset. Therefore, we offer a new perspective by exploring the concept of generalization error from a data-centered perspective. Initially, we establish the following lemmas, which are essential to derive the upper bound on generalization error.
Lemma 2.
Given any distributions $q,\bar{q}$ on the space $\mathcal{X}$ and a function $f:\mathcal{X}\to[0,L]$, the following inequality holds: $D(q\|\bar{q})\geq\frac{2}{L^{2}}(\mathbb{E}_{q}f-\mathbb{E}_{\bar{q}}f)^{2}$.
Proof.
Begin with the following calculation:
$|\mathbb{E}_{q}$ $f-\mathbb{E}_{\bar{q}}f|=|\sum_{x\in\mathcal{X}}(q(x)-\bar{q}(x))f(x)|$ (15)
$=|\sum_{x\in\mathcal{X}}(\bar{q}(x)-q(x))(f(x)-\frac{L}{2})+\frac{L}{2}(\bar{q}(x)-q(x))|$
$\leq\sum_{x\in\mathcal{X}}|\bar{q}(x)-q(x)||f(x)-\frac{L}{2}|$
$\leq\frac{L}{2}\|q-\bar{q}\|_{1}$
Pinsker’s inequality[56, 57] states that:
$D(q\|\bar{q})\geq\frac{1}{2}\|q-\bar{q}\|_{1}^{2}.$ (16)
Combine the two inequalities (15) and (16), we obtain:
$D(q\|\bar{q})\geq\frac{2}{L^{2}}|\mathbb{E}_{q}f-\mathbb{E}_{\bar{q}}f|^{2}$ (17)
To establish bound on the generalization error, we introduce the following theorem.
Theorem 1 (Bound on generalization error).
$\Pr(\mathrm{gen}(\ell,f_{\theta})\geq\epsilon)\leq\frac{L^{2}\mathbb{E}_{\bar{q}|q}D_{KL}(q\|\bar{q})}{2\epsilon^{2}}$ (18)
where $L=\|\ell(y,f_{\theta}(x))\|_{\infty}$, $nq\sim\operatorname{Multinomial}(n,\bar{q})$.
Proof.
Since $q$ is the empirical PMF of $(X,Y)$ based on the training set $S^{n}$, Lemma 2 implies that $D_{KL}(q\|\bar{q})\geq\frac{2}{L^{2}}\mathrm{gen}(\ell,f_{\theta})^{2}$, where $L=\|\ell(y,f_{\theta}(x))\|_{\infty}$. This leads to:
$\Pr(D_{KL}(q\|\bar{q})\geq t)$ $\geq\Pr(\frac{2}{L^{2}}\mathrm{gen}(\ell,f_{\theta})^{2}\geq t)$ (19)
$=\Pr(\mathrm{gen}(\ell,f_{\theta})^{2}\geq L^{2}t/2)$
$=\Pr(\mathrm{gen}(\ell,f_{\theta})\geq L\sqrt{t/2}).$
Let $\epsilon=L\sqrt{t/2}$, because $\mathrm{gen}(\ell,f_{\theta}),D_{KL}(q\|\bar{q})$ are functions of the random variable $\bar{q}$, applying Lemma 1, we obtain:
$\Pr(\mathrm{gen}(\ell,f_{\theta})\geq\epsilon)$ $\leq\Pr(D_{KL}(q\|\bar{q})\geq\frac{2\epsilon^{2}}{L^{2}})$ (20)
$\leq\frac{L^{2}\mathbb{E}_{\bar{q}|q}D_{KL}(q\|\bar{q})}{2\epsilon^{2}}.$
As shown in Theorem 1, the generalization error is influenced not only by the maximum loss $\ell(\ell,f_{\theta})$ but also by $\frac{L^{2}\mathbb{E}_{\bar{q}|q}D_{KL}(q\|\bar{q})}{2\epsilon^{2}}$. Consequently, we define the complexity of the learning task (i.e., the complexity of dataset) as
$C(q)=\frac{\mathbb{E}_{\bar{q}|q}D_{KL}(q\|\bar{q})}{2},$ (21)
where $nq\sim\operatorname{Multinomial}(n,\bar{q})$. To estimate complexity efficiently, we present the following propositions.
Proposition 2.
  1. 1.
    As the sampling size escalates, the uncertainty tends asymptotically towards zero:
    $\lim_{n\to\infty}C(q)=0$ (22)
  2. 2.
    If the sample data is indifferent and all existing constraints are satisfied, a more uniform $q$ implies less uncertainty.
Proof.
According to the Bayesian theorem and Sanov’s theorem [55], we have
[…]
Combining the above equations, we obtain the following limits as $n$ approaches infinity:
$\lim_{n\to\infty}\Pr(\bar{q}=q|q)$ $=1$ (24)
$\lim_{n\to\infty}\mathbb{E}_{\bar{q}|q}D_{KL}(q\|\bar{q})$ $=0.$
First, let’s find the lower bound of $C(q)$:
$C(q)$ $=-H(q)-\mathbb{E}_{\bar{q}|q}\mathbb{E}_{Z\sim q}\log\bar{q}(z)$ (25)
$=-H(q)-\mathbb{E}_{Z\sim q}\mathbb{E}_{\bar{q}|q}\log\bar{q}(z)$
$\geq-H(q)-\mathbb{E}_{Z\sim q}\log\mathbb{E}_{\bar{q}|q}\bar{q}(z)$
where the final inequality is derived from the Jensen’s inequality, as $-\log(\cdot)$ is a convex function. Given that $\sum_{Z}\mathbb{E}_{\bar{q}|q}\bar{q}(z)=1$, therefore we have
$C(q)\geq D_{KL}(q\|\mathbb{E}_{\bar{q}|q}\bar{q}).$ (26)
Then the upper bound of $C(q)$ is derived as follows:
$C(q)$ $=-H(q)-\mathbb{E}_{\bar{q}|q}\mathbb{E}_{Z\sim q}\log\bar{q}(z)$ (27)
$=-H(q)+\mathbb{E}_{\bar{q}|q}\mathbb{E}_{Z\sim q}[1+\frac{1-\bar{q}(z)}{\bar{q}(z)}]$
$\leq-H(q)+\mathbb{E}_{\bar{q}|q}\mathbb{E}_{Z\sim q}\frac{1-\bar{q}(z)}{\bar{q}(z)}$
where the final inequality is base on $\log(1+x)\leq x$. Combine the formula (26) and (27), we have
$D_{KL}(q\|\mathbb{E}_{\bar{q}|q}\bar{q})$ $\leq C(q)$ (28)
$\leq-H(q)+\mathbb{E}_{\bar{q}|q}\mathbb{E}_{Z\sim q}\frac{1-\bar{q}(z)}{\bar{q}(z)}$
$\leq H(q)+\mathbb{E}_{\bar{q}|q}\sum_{Z}\frac{q_{\max}}{\bar{q}(z)}-1.$
Because $\bar{q}$ is an arbitrary distribution on $\mathcal{Z}$ , when $n$ is insufficient large, the conditional probability $\Pr(\bar{q}=q|q)$ fails to approximate to 1, causing $\mathbb{E}_{\bar{q}|q}\bar{q}$ to be a relatively uniform. Consequently, a more uniform $q$ often implies a smaller lower $D_{KL}(q\|\mathbb{E}_{\bar{q}|q}\bar{q})$, i.e., smaller bound of $C(q)$. The fact that $q_{\max}$ influences the upper bound also highlights the advantage of a uniform $q$ in minimizing the $C(q)$. Thus, under the condition of satisfying existing constraints, a more uniform $q$ implies lower uncertainty. ∎
According to Proposition 2, given the assumption of i.i.d. sampling, the upper limit of complexity is directly related to the smoothness of $q$ and the sample size. This conclusion is consistent with the law of large numbers: more samples lead to reduced bias from random sampling. The principle of maximum entropy posits that the most probable distribution subject to specified constraints should be the maximum entropy distribution. Furthermore, $H(q)$ serves as an indicator of the uniformity of $q$, thus establishing a degree of equivalence between the negative Shannon entropy $-H(q)$ and the uncertainty $C(q)$.
As posited by Theorem 1, the complexity $C(q)$ serves as an indicator of the discrepancy between the training dataset and the entire population, reflecting the reliability of the training data. Consequently, we harness this complexity to inform the selection of hyperparameters for regularization term. Intuitively, an increase in complexity of classification tasks necessitates a heightened focus on the regularization loss. Furthermore, when the model perfectly aligns with the training data $S^{n}$ (i.e., $p=q$), the complexity can serve as an indicator of confidence in the model’s predictions. In order to compare our proposed complexity with the classical complexity of hypothesis space, we introduce the expected generalization error, defined as $2\mathbb{E}\mathrm{gen}(\ell,f_{\theta})^{2}/L^{2}$. According to Theorem 1, this expected error is bounded by $C(q)$, that is, $2\mathbb{E}\mathrm{gen}(\ell,f_{\theta})^{2}/L^{2}\leq C(q)$.
The interrelation between expected generalization error and complexity is summarized in the curve shown in Figure 2 (b). For comparison, the double descent risk curve is shown in Figure 2 (a).
(a) The double descent risk curve, which incorporates the U-shaped risk curve [13].
(b) Curves for expected generalization error and complexity.
Figure 2: The effect of the capacity of the hypothesis space and task complexity on the generalization error.
The double descent risk curve, which incorporates the U-shaped risk curve (i.e., the “classical” regime) together with the observed behavior from using high capacity function classes (i.e., the “modern” interpolating regime), separated by the interpolation threshold. The figure posits that the capacity of the hypothesis space, as quantified by classical complexity, does not exhibit a substantial linear relationship with generalization error across the entire domain. In contrast, our proposed complexity metric is able to tightly manage the expected generalization error. In summary, we demonstrate that the generalization error can be controlled by the complexity of the classification task, and conclude that the number of training samples as well as the uniformity of the distribution jointly determine the complexity.
5 Fitting error and influencing factors
In this section, we aim to bound the fitting error, and examine the influence of the model’s architecture on it. To do this, we begin by introducing several key terms used to derive bound of fitting error as follows:
$\ell(f_{\theta}(x))=(\ell(f_{\theta}(x),y_{1}),\cdots,\ell(f_{\theta}(x),y_{|\mathcal{Y}|}))^{\top},$ (29)
$F(\theta,q_{Y|x})=\frac{\|\nabla_{\theta}D_{KL}(q_{Y|x}\|p_{Y|x})\|_{2}^{2}}{\|\nabla_{\theta}f_{\theta}(x)\|_{F}^{2}},$
$F(\theta_{j},q_{Y|x})=\frac{\|\nabla_{\theta_{j}}D_{KL}(q_{Y|x}\|p_{Y|x})\|_{2}^{2}}{\|\nabla_{\theta}f_{\theta}(x)\|_{2}^{2}},$
$G(\theta,q_{Y|x})=\frac{\sum_{i=1}^{m}\|(q_{Y|x}-p_{Y|x})\times\nabla_{\theta_{i}}f_{\theta}(x)\|_{2}^{2}}{\|\nabla_{\theta}f_{\theta}(x)\|_{F}^{2}},$
$G(\theta_{j},q_{Y|x})=\frac{\|(q_{Y|x}-p_{Y|x})\times\nabla_{\theta_{j}}f_{\theta}(x)\|_{2}^{2}}{\|\nabla_{\theta_{j}}f_{\theta}(x)\|_{2}^{2}},$
where $nq\sim\operatorname{Multinomial}(n,\bar{q})$, $m$ represents the number of parameters in the model, and $\theta_{j}$ denotes the $j$-th parameter of the model. First, we present two lemmas, essential for establishing an upper bound on the fitter error.
Lemma 3.
Given $X,Y$ are two random variables, with finite second moments, i.e., $\mathbb{E}[X^{2}]<\infty,\mathbb{E}[Y^{2}]<\infty$, the following inequality holds:
$(\mathbb{E}[\langle X,Y\rangle])^{2}\leq\mathbb{E}[\|X\|^{2}]\mathbb{E}[\|Y\|^{2}].$ (30)
Proof.
Consider the function $f(\alpha)=\mathbb{E}[\|\alpha X-Y\|^{2}]$ , which upon expansion yields quadratic form:
$f(\alpha)=\alpha^{2}~{}\mathbb{E}[\|X\|^{2}]-2~{}\alpha~{}\mathbb{E}[~{}|\langle X,Y\rangle|~{}]+\mathbb{E}[\|Y\|^{2}]\geq 0.$ (31)
The discriminant of this quadratic function $\Delta=b^{2}-4ac$, is given by:
$\Delta=b^{2}-4ac=\mathbb{E}[~{}2~{}|\langle X,Y\rangle|~{}]^{2}-4~{}\mathbb{E}[\|X\|^{2}]~{}\mathbb{E}[\|Y\|^{2}].$ (32)
For $f$ has at most one root, the discriminant must be non-positive, i.e., $\Delta\leq 0$, and the equality holds iif $\alpha X=Y$. ∎
Lemma 4.
$\|Ax\|_{2}^{2}=\|A\|^{2}_{F}\|x\|_{2}^{2}-\sum_{k=1}^{m}\|a_{k}\times x\|_{2}^{2}$, where $A$ is a $m\times n$ matrix, $x\in\mathbb{R}^{n}$, $\|\cdot\|_{2}$ is the Euclidean norm, and $\|\cdot\|_{F}$ is the Frobenius norm. and $\|a_{k}\times x\|_{2}^{2}=\frac{1}{2}\sum_{i\mathop{=}1}^{n}\sum_{j\mathop{=}1}^{n}(a_{ki}x_{j}-a_{kj}x_{i})^{2}$.
Proof.
Because the following equation always holds
[…]
we obtain
$\sum_{k=1}^{m}$ $\Big{(}\sum_{j=1}^{n}a_{kj}x_{j}\Big{)}^{2}=$ (34)
$\sum_{k=1}^{m}\sum_{j=1}^{n}a_{kj}^{2}\sum_{j=1}^{n}x_{j}^{2}-\sum_{k=1}^{m}\|a_{k}\times x\|_{2}^{2}.$
For
$\|Ax\|^{2}_{2}=\sum_{k=1}^{m}(\sum_{j=1}^{n}a_{kj}x_{j})^{2},$ (35)
then we obtain
$\|Ax\|_{2}^{2}=\|A\|^{2}_{F}\|x\|_{2}^{2}-\sum_{k=1}^{m}\|a_{k}\times x\|_{2}^{2}.$ (36)
The subsequent theorem provides an upper bound on the fitting error.
Theorem 3 (Bound on fitting error).
$\mathrm{fit_{n}}(\ell,f_{\theta})$ $\leq\sqrt{\mathbb{E}_{X}[F(\theta,q_{Y|x})]+\mathbb{E}_{X}[G(\theta,q_{Y|x})]}$ (37)
$=\sqrt{\mathbb{E}_{X}[F(\theta_{j},q_{Y|x})]+\mathbb{E}_{X}[G(\theta_{j},q_{Y|x})]}$
where $X\sim q_{X}$.
Proof.
By inserting
$p_{Y|x}(y)=\frac{e^{f_{\theta}(x)}}{Z_{\theta}(x)},$ (38)
into $D_{KL}(q_{Y|x}\|p_{Y|x})$, we obtain
$D_{KL}(q_{Y|x}\|p_{Y|x})=q_{Y|x}^{\top}f_{\theta}(x)-\log Z_{\theta}(x)-H(q_{Y|x}).$ (39)
Taking the gradient of function $D_{KL}(q_{Y|x}\|p_{Y|x})$ with respect to $\theta$, we have:
$\nabla_{\theta}D_{KL}(q_{Y|x}\|p_{Y|x})=\nabla_{\theta}f_{\theta}(x)^{\top}(q_{Y|x}-p_{Y|x}).$ (40)
Applying Lemma 4 , we derive:
$\|\nabla_{\theta}$ $D_{KL}(q_{Y|x}\|p_{Y|x})\|_{2}^{2}=\|q_{Y|x}-p_{Y|x}\|_{2}^{2}\|\nabla_{\theta}f_{\theta}(x)\|_{F}^{2}$ (41)
$-\sum_{i=1}^{|\mathcal{Y}|}\|(q_{Y|x}-p_{Y|x})\times\nabla_{\theta_{i}}f_{\theta}(x)\|_{2}^{2},$
where $m$ is the number of parameters. Extracting $\|q_{Y|x}-p_{Y|x}\|_{2}^{2}$ from this expression, we get:
$\|q_{Y|x}-$ $p_{Y|x}\|_{2}^{2}=\frac{\|\nabla_{\theta}D_{KL}(q_{Y|x}\|p_{Y|x})\|_{2}^{2}}{\|\nabla_{\theta}f_{\theta}(x)\|_{F}^{2}}$ (42)
$+\frac{\sum_{i=1}^{m}\|(q_{Y|x}-p_{Y|x})\times\nabla_{\theta_{i}}f_{\theta}(x)\|_{2}^{2}}{\|\nabla_{\theta}f_{\theta}(x)\|_{F}^{2}}.$
Repeating the process for the derivative with respect to single parameter $\theta_{j}$, we obtain similar conclusion
$\|q_{Y|x}-p_{Y|x}\|_{2}^{2}$ $=\frac{\|\nabla_{\theta_{j}}D_{KL}(q_{Y|x}\|p_{Y|x})\|_{2}^{2}}{\|\nabla_{\theta_{j}}f_{\theta}(x)\|_{F}^{2}}$ (43)
$+\frac{\|(q_{Y|x}-p_{Y|x})\times\nabla_{\theta_{j}}f_{\theta}(x)\|_{2}^{2}}{\|\nabla_{\theta_{j}}f_{\theta}(x)\|_{2}^{2}},$
According to the definitions of $\ell(f_{\theta}(x)),F(\theta,q_{Y|x},F(\theta_{j},q_{Y|x}),G(\theta,q_{Y|x})),G(\theta_{j},q_{Y|x})$, we have
$\|q_{Y|x}-p_{Y|x}\|_{2}^{2}$ $=F(\theta,q_{Y|x})+G(\theta,q_{Y|x})$ (44)
$=F(\theta_{j},q_{Y|x})+G(\theta_{j},q_{Y|x})$
where $j\in[1,m]$, $m$ is the number of parameters.
Using Lemma 3 and Equation (44), we can bound the expected absolute difference as follows:
$\mathbb{E}_{X}$ $|p_{Y|x}^{\top}\ell(f_{\theta}(x))-q_{Y|x}^{\top}\ell(f_{\theta}(x))|$ (45)
$\leq\sqrt{\mathbb{E}_{X}[\|p_{Y|x}-q_{Y|x}\|^{2}_{2}]\mathbb{E}_{x}[\|\ell(f_{\theta}(x))\|^{2}_{2}]}$
$=\sqrt{\mathbb{E}_{X}[F(\theta,q_{Y|x})+G(\theta,q_{Y|x})]\mathbb{E}_{x}[\|\ell(f_{\theta}(x))\|^{2}_{2}]}$
$=\sqrt{\mathbb{E}_{X}[F(\theta_{j},q_{Y|x})+G(\theta_{j},q_{Y|x})]\mathbb{E}_{x}[\|\ell(f_{\theta}(x))\|^{2}_{2}]}.$
Since $p(x,y)=q_{X}(x)p_{Y|x}(y)$, we have:
$\mathbb{E}_{(X,Y)\sim p}\ell(f_{\theta}(x),y)$ $=\frac{1}{n}\sum_{i=1}^{n}\ell([f_{\theta}(x)]_{i},y_{i})$ (46)
$=\mathbb{E}_{X\sim q_{X}}p_{Y|x}^{\top}\ell(f_{\theta}(x)).$
This leads to:
$|\mathbb{E}_{(X,Y)\sim p}\ell(f_{\theta}(x),y)-\mathbb{E}_{(X,Y)\sim q}\ell(f_{\theta}(x),y)|$ (47)
$\leq\sqrt{\mathbb{E}_{X}[F(\theta,q_{Y|x})+G(\theta,q_{Y|x})]\mathbb{E}_{X}[\|\ell(f_{\theta}(x))\|^{2}_{2}]}$
$=\sqrt{\mathbb{E}_{X}[F(\theta_{j},q_{Y|x})+G(\theta_{j},q_{Y|x})]\mathbb{E}_{X}[\|\ell(f_{\theta}(x))\|^{2}_{2}]}$
Divide both sides of the inequality by $\sqrt{\mathbb{E}_{X}[\|\ell(f_{\theta}(x))\|^{2}_{2}]}$, we obtain
$\frac{\mathrm{fit}(\ell,f_{\theta})}{\sqrt{\mathbb{E}_{X}[\|\ell(f_{\theta}(x))\|^{2}_{2}]}}$ $\leq\sqrt{\mathbb{E}_{X}[F(\theta,q_{Y|x})+G(\theta,q_{Y|x})]}$ (48)
$=\sqrt{\mathbb{E}_{X}[F(\theta_{j},q_{Y|x})+G(\theta_{j},q_{Y|x})]}.$
The term $\|\nabla_{\theta}D_{KL}(q_{Y|x}\|p_{Y|x})\|_{2}^{2}$ represents the gradient of the KL divergence between $q_{Y|x}$ and $p_{Y|x}$ with respect to $\theta$. Since the function $\ell(f_{\theta}(x),y)$ is smooth and differentiable with respect to $\theta$, gradient-based algorithms iterate the parameters toward the stationary points where $\mathbb{E}_{X,Y}\nabla_{\theta}\ell(f_{\theta}(x),y)=0$. Consequently, if $R_{SEC}(f_{\theta},q)$ is a component of the total loss function for model training, gradient-based algorithms can effectively reduce $\|\nabla_{\theta}D_{KL}(q_{Y|x}\|p_{Y|x})\|_{2}^{2}$ for each input feature $x$.
The above theorem highlights that even when gradient-based algorithms fail to reach the global optimal solution due to the non-convexity of loss function, controlling the upper bound of normalized fitting error allows for an approximation of the global optima. Notably, cross-entropy, a prevalent choice in classification tasks, is equivalent to KL divergence when computing the gradient with respect to $\theta$. Consequently, employing cross-entropy as the loss function enables gradient-based algorithms can minimize the upper bound on the normalized fitting error.
Since
$\|\nabla_{\theta}D_{KL}(q_{Y|x}\|p_{Y|x})\|_{2}^{2}=\|\nabla_{\theta}f_{\theta}(x)^{\top}(q_{Y|x}-p_{Y|x})\|_{2}^{2},$ (49)
this equation leads to the inequality:
$\|\nabla_{\theta}D_{KL}(q_{Y|x}\|p_{Y|x})\|_{2}^{2}$ $\leq\lambda_{\max}\|(q_{Y|x}-p_{Y|x})\|_{2}^{2},$ (50)
where $\lambda_{\max}$ is the largest eigenvalue of eNTK, i.e., $\nabla_{\theta}f_{\theta}(x)^{\top}\nabla_{\theta}f_{\theta}(x)$. According to the definition $F(\theta,q_{Y|x})$, we obtain the following bound:
$F(\theta,q_{Y|x})\leq\frac{\lambda_{\max}\|q_{Y|x}-p_{Y|x}\|_{2}^{2}}{\mathrm{trace}(H(f_{\theta}))}.$ (51)
Hence, minimizing $\lambda_{\max}$ is equivalent to minimizing $F(\theta,q_{Y|x})$, which gives rise to the following corollary.
Corollary 1 (Minimum largest eigenvalue).
Reducing the largest eigenvalue of $H(f_{\theta}(x))$ is advantageous for controlling fitting errors.
In the flat minima theory, it is posited that flat minima generalize better than sharp (non-flat) minima, and the largest eigenvalue of the Hessian of the loss with respect to the model parameters on the training set serves as a metric for flatness. Similar to flat minima but with a twist, we have demonstrated that decreasing the largest eigenvalue of eNTK helps in reducing fitting errors. We now proceed to establish the equivalence between the two concepts, thereby indirectly substantiating the efficacy and validity of our proposed bound. Prior to presenting the theorem, we provide following lemmas, which are instrumental in proving the theorem.
Lemma 5.
Let $A=(a_{ij})_{m\times m}$ be a symmetric matrix. Then, we have
$a_{ii}\leq\lambda_{\max}(A)\leq\mathrm{trance}(A)\leq m\lambda_{\max}(A),\forall i\leq m.$ (52)
Proof.
Given $A=(a_{ij})_{m\times m}$ be a symmetric matrix, then we have $\lambda_{\min}(A)\geq 0$. Since the sum of eigenvalue is equal to $\mathrm{trance}(A)$, thus we have
$\lambda_{\max}(A)\leq\mathrm{trance}(A)\leq m\lambda_{\max}(A).$ (53)
By Courant-Fisher min-max theorem, we have
$\lambda_{\max}(A)=\max_{x\in\mathbb{R}^{m},\|x\|=1}x^{\top}Ax.$ (54)
Let $e_{i}$ be a standard basis (“one-hot”) vector, then we have
$a_{ii}=e_{i}^{\top}Ae_{i}\leq\lambda_{\max}(A)$ (55)
Lemma 6.
Let $A,B$ be non-negative definite matrices. Then we have
$\lambda_{\min}(A+B)$ $\geq\lambda_{\min}(A)+\lambda_{\min}(B)$ (56)
$\lambda_{\max}(A+B)$ $\leq\lambda_{\max}(A)+\lambda_{\max}(B).$
Proof.
Because
$x^{\top}Ax\geq\lambda_{\min}(A)x^{\top}x,$ (57)
$x^{\top}(A-\lambda_{\min}(A)I)x\geq 0,$
thus $A-\lambda_{\min}(A)I$ is non-negative definite matrix. Therefore, we have
$A-\lambda_{\min}(A)I\succeq 0$ (58)
$B-\lambda_{\min}(B)I\succeq 0$
$A-\lambda_{\min}(A)I+B-\lambda_{\min}(B)I\succeq 0.$
Let $(A+B)x^{\prime}=\lambda_{\min}(A+B)x^{\prime}$, then we have
$(A-\lambda_{\min}(A)I+B$ $-\lambda_{\min}(B)I)x^{\prime}=(\lambda_{\min}(A+B)$ (59)
$-\lambda_{\min}(A)-\lambda_{\min}(B))x^{\prime}.$
Therefore $\lambda_{\min}(A+B)-\lambda_{\min}(A)-\lambda_{\min}(B)$ is the eigenvalue of $A-\lambda_{\min}(A)I+B-\lambda_{\min}(B)I$. Because $A-\lambda_{\min}(A)I+B-\lambda_{\min}(B)I$ is non-negative definite matrix, thus we have $\lambda_{\min}(A+B)-\lambda_{\min}(A)-\lambda_{\min}(B)\geq 0$.
Because $x^{\top}A\leq\lambda_{\max}(A)x^{\top}x$, then $x^{\top}(A-\lambda_{\max}(A)I)x\leq 0$, thus $\lambda_{\max}(A)I-A$ is non-negative definite matrix. Similarly, we can obtain $\lambda_{\max}(A+B)-\lambda_{\max}(A)-\lambda_{\max}(B)\leq 0$. ∎
Proposition 4.
If $R_{ERF}(f_{\theta},q)$ is employed as loss function, and the following equations holds
$\sum_{i=1}^{|\mathcal{Y}|}p_{Y|x}(y_{i})\nabla_{\theta_{i}}f_{\theta}(x)_{i}=0,$ (60)
$\mathbb{E}_{X\sim q_{X}}\sum_{i=1}^{|\mathcal{Y}|}p_{Y|x}(y_{i})\frac{\nabla_{\theta}^{2}p_{Y|x}(y_{i})}{p_{Y|x}(y_{i})}=0,$
after training, then we have:
$\lambda_{\max}(H)\leq\max_{x}\lambda_{\max}(H(f_{\theta}(x))),$ (61)
where $H$ is the Hessian for $R_{ERF}(f_{\theta},q)$.
Proof.
According to the definition, we have $p_{Y|x}(y_{i})=\frac{e^{[f_{\theta}(x)]_{i}}}{Z_{\theta}(x)}$, $Z_{\theta}(x)$ is the partition function, defined as $Z_{\theta}(x)=\sum_{j=1}^{|\mathcal{Y}|}e^{f_{\theta}(x)_{j}}$. Consequently, the gradient of the log-probability with respect to the model parameter $\theta$ decomposes as follows:
$\nabla_{\theta}\log p_{Y|x}(y)=\nabla_{\theta}[f_{\theta}(x)]_{j}-\nabla_{\theta}\log Z_{\theta}(x),$ (62)
The second gradient term, $\nabla_{\theta}\log Z_{\theta}(x)$ can be rewritten as an expected value, as follows:
$\nabla_{\theta}\log Z_{\theta}(x)$ $=Z_{\theta}(x)^{-1}\nabla_{\theta}\sum_{j=1}^{|\mathcal{Y}|}e^{f_{\theta}(y)_{j}}$ (63)
$=Z_{\theta}(x)^{-1}\sum_{j=1}^{|\mathcal{Y}|}e^{f_{\theta}(y)_{j}}\nabla_{\theta}f_{\theta}(y)_{j}$
$=\sum_{j=1}^{|\mathcal{Y}|}q_{Y|x}(y_{j})\nabla_{\theta}[f_{\theta}(x)]_{j}.$
According to the above equality, we have
$\sum_{i=1}^{|\mathcal{Y}|}q_{Y|x}(y_{i})$$\left\{\nabla_{\theta}\log p_{Y|x}(y)\right\}\left\{\nabla_{\theta}\log p_{Y|x}(y)\right\}^{\top}$ (64)
$=\sum_{i=1}^{|\mathcal{Y}|}q_{Y|x}(y_{i})\{\nabla_{\theta}[f_{\theta}(x)]_{i}-\nabla_{\theta}\log Z_{\theta}(x)\}$
$\{\nabla_{\theta}[f_{\theta}(x)]_{i}-\nabla_{\theta}\log Z_{\theta}(x)\}^{\top}$
$=\sum_{i=1}^{|\mathcal{Y}|}q_{Y|x}(y_{i})\{\nabla_{\theta}[f_{\theta}(x)]_{i}\nabla_{\theta}[f_{\theta}(x)]_{i}^{\top}\}$
$-2\nabla_{\theta}\log Z_{\theta}(x)\mathbb{E}_{Y\sim q_{Y|x}}\nabla_{\theta}[f_{\theta}(x)]_{i}$
$+\nabla_{\theta}\log Z_{\theta}(x)\nabla_{\theta}\log Z_{\theta}(x)^{\top}$
$=\sum_{i=1}^{|\mathcal{Y}|}q_{Y|x}(y_{i})\{\nabla_{\theta}[f_{\theta}(x)]_{i}\nabla_{\theta}[f_{\theta}(x)]_{i}^{\top}\}+F_{x},$
where
$F_{x}$ $=-2\nabla_{\theta}\log Z_{\theta}(x)\sum_{i=1}^{|\mathcal{Y}|}q_{Y|x}(y_{i})\nabla_{\theta}[f_{\theta}(x)]_{i}$ (65)
$+\nabla_{\theta}\log Z_{\theta}(x)\nabla_{\theta}\log Z_{\theta}(x)^{\top}.$
The Hessian for $D_{KL}(q_{Y|x},p_{Y|x})$ is equal to compute with the second derivative of negative log-likelihood $-\log p_{Y|x}(y)$, which can be expressed as:
$\nabla^{2}_{\theta}D_{KL}(q_{Y|x},p_{Y|x})=\sum_{i=1}^{|\mathcal{Y}|}q_{Y|x}(y_{i})[-\nabla^{2}_{\theta}\log p_{Y|x}(y_{i})].$ (66)
Next, we expand the second derivative of it as follows:
$-\nabla_{\theta}^{2}$ $\log p_{Y|x}(y)=-\nabla_{\theta}\left[\frac{\nabla_{\theta}p_{Y|x}(y)}{p_{Y|x}(y)}\right]$ (67)
$=-\frac{p_{Y|x}(y)\nabla_{\theta}^{2}p_{Y|x}(y)-\left\{\nabla_{\theta}p_{Y|x}(y)\right\}\left\{\nabla_{\theta}p_{Y|x}(y)\right\}^{\top}}{p_{Y|x}(y)^{2}}$
$=-\frac{\nabla_{\theta}^{2}p_{Y|x}(y)}{p_{Y|x}(y)}+\left\{\frac{\nabla_{\theta}p_{Y|x}(y)}{p_{Y|x}(y)}\right\}\left\{\frac{\nabla_{\theta}p_{Y|x}(y)}{p_{Y|x}(y)}\right\}^{\top}$
$=-\frac{\nabla_{\theta}^{2}p_{Y|x}(y)}{p_{Y|x}(y)}+\left\{\nabla_{\theta}\log p_{Y|x}(y)\right\}\left\{\nabla_{\theta}\log p_{Y|x}(y)\right\}^{\top}.$
Then, we apply this expanded second derivative to (66) which can be restated as follows:
$\nabla^{2}_{\theta}$ $D_{KL}(q_{Y|x},p_{Y|x})=\sum_{i=1}^{|\mathcal{Y}|}q_{Y|x}(y_{i})[-\frac{\nabla_{\theta}^{2}p_{Y|x}(y_{i})}{p_{Y|x}(y_{i})}$ (68)
+$+\left\{\nabla_{\theta}\log p_{Y|x}(y_{i})\right\}\left\{\nabla_{\theta}\log p_{Y|x}(y_{i})\right\}^{\top}]$
$=C_{x}+B_{x}+F_{x}.$
where
$B_{x}$ $=\sum_{i=1}^{|\mathcal{Y}|}q_{Y|x}(y_{i})\nabla_{\theta}[f_{\theta}(x)]_{i}\nabla_{\theta}[f_{\theta}(x)]_{i}^{\top},$ (69)
$C_{x}$ $=\sum_{i=1}^{|\mathcal{Y}|}q_{Y|x}(y_{i})\left[-\frac{\nabla_{\theta}^{2}p_{Y|x}(y_{i})}{p_{Y|x}(y_{i})}\right].$
Since $\nabla_{\theta}[f_{\theta}(x)]_{i}\nabla_{\theta}[f_{\theta}(x)]_{i}^{\top}$ is matrix of rank 1, its eigenvalues are
$\lambda_{\max}(\nabla_{\theta}[f_{\theta}(x)]_{i}\nabla_{\theta}[f_{\theta}(x)]_{i}^{\top})$ $=\nabla_{\theta}[f_{\theta}(x)]_{i}^{\top}\nabla_{\theta}[f_{\theta}(x)]_{i}$ (70)
$\lambda_{\min}(\nabla_{\theta}[f_{\theta}(x)]_{i}\nabla_{\theta}[f_{\theta}(x)]_{i}^{\top})$ $=0$
According to the lemma 6, we can bounded the extreme eigenvalues of $B_{x}$ as follows:
$\lambda_{\max}(B_{x})$ $\leq\sum_{i=1}^{|\mathcal{Y}|}q_{Y|x}(y_{i})\lambda_{\max}(\nabla_{\theta}[f_{\theta}(x)]_{i}\nabla_{\theta}[f_{\theta}(x)]_{i}^{\top})$ (71)
$\leq\max_{y}\lambda_{\max}(\nabla_{\theta}[f_{\theta}(x)]_{i}\nabla_{\theta}[f_{\theta}(x)]_{i}^{\top})$
$=\max_{y}\nabla_{\theta}[f_{\theta}(x)]_{i}^{\top}\nabla_{\theta}[f_{\theta}(x)]_{i}$
$\lambda_{\min}(B_{x})$ $\geq\sum_{i=1}^{|\mathcal{Y}|}q_{Y|x}(y_{i})\lambda_{\min}(\nabla_{\theta}[f_{\theta}(x)]_{i}\nabla_{\theta}[f_{\theta}(x)]_{i}^{\top})$
$\geq\min_{y}\lambda_{\min}(\nabla_{\theta}[f_{\theta}(x)]_{i}\nabla_{\theta}[f_{\theta}(x)]_{i}^{\top})$
$=0.$
Given that $H(f_{\theta}(x))=\nabla_{\theta}f_{\theta}(x)^{\top}\nabla_{\theta}f_{\theta}(x)$ which is a symmetric matrix, according to the inequality (71) and lemma5 , we have:
$\lambda_{\max}(B_{x})\leq\max_{y}\nabla_{\theta}[f_{\theta}(x)]_{i}^{\top}\nabla_{\theta}[f_{\theta}(x)]_{i}$ $\leq\lambda_{\max}(H(f_{\theta}(x))).$ (72)
The Hessian for $\mathbb{E}_{X\sim q_{X}}D_{KL}(q_{Y|x},p_{Y|x})$, can be denoted as:
$H=\mathbb{E}_{X\sim q_{X}}B_{x}+\mathbb{E}_{X\sim q_{X}}C_{x}+\mathbb{E}_{X\sim q_{X}}F_{x}.$ (73)
Since
$\sum_{i=1}^{|\mathcal{Y}|}q_{Y|x}(y_{i})\nabla_{\theta}[f_{\theta}(x)]_{i}=0,$ (74)
$\mathbb{E}_{X\sim q_{X}}\sum_{i=1}^{|\mathcal{Y}|}q_{Y|x}(y_{i})\frac{\nabla_{\theta}^{2}p_{Y|x}(y_{i})}{p_{Y|x}(y_{i})}=0,$
and since $H=\mathbb{E}_{X\sim q_{X}}B_{x}$, we deduce
$\lambda_{\max}(H)\leq\max_{x}\lambda_{\max}(B_{x})\leq\max_{x}\lambda_{\max}(H(f_{\theta}(x)))$ (75)
As shown in proposition 4, minimizing $\lambda_{\max}(H(f_{\theta}(x))),\forall x\in\mathcal{X}$, is equivalent to minimizing $\lambda_{\max}(\nabla^{2}_{\theta}R_{SEC}(f_{\theta},q))$ and our proposed bound on fitting error provide an alternative theoretical explanation for the flat minima theory.
Based on the definition of $G(\theta_{i},q_{Y|x})$, we can express it as:
$G(\theta_{i},q_{Y|x})=\|q_{Y|x}-p_{Y|x}\|^{2}_{2}\sin^{2}\alpha(\theta_{i},x),$ (76)
where $\alpha(\theta_{i},x)$ denotes the angle between vectors $(q_{Y|x}-p_{Y|x})$ and $\nabla_{\theta_{i}}f_{\theta}(x)$. Gradient-based algorithms do not directly influence $\alpha(\theta_{i},x)$ during the optimization of the empirical risk. Therefore, assuming $\alpha(\theta_{i},x)$ behaves as a random variable after training is reasonable. Considering a model with $m$ parameters, $\{\alpha(\theta_{1},x),\cdots,\alpha(\theta_{m},x)\}$ can be regarded as $m$ random samples from a distribution. We define $G_{M}(\theta_{1}^{m},q_{Y|x})$ as the minimum of $\{\mathbb{E}_{X}G(\theta_{1},q_{Y|x}),\cdots,\mathbb{E}_{X}G(\theta_{m},q_{Y|x})\}$:
$G_{M}(\theta_{1}^{m},q_{Y|x})=\min\{\mathbb{E}_{X}G(\theta_{1},q_{Y|x}),\cdots,\mathbb{E}_{X}G(\theta_{m},q_{Y|x})\},$ (77)
we can derive the probability that $G_{M}$ does not exceed a constant $c$ as follows:
$\Pr(G_{M}$ $(\theta_{1}^{m},q_{Y|x})\leq c)$ (78)
$=1-\Pr(\cap_{i=1}^{m}\{\mathbb{E}_{X}G(\theta_{i},q_{Y|x})\geq c\})$
$\geq 1-\Pr(\cap_{i=1}^{m-1}\{\mathbb{E}_{X}G(\theta_{i},q_{Y|x})\geq c\})$
$=\Pr(G_{M}(\theta_{1}^{m-1},q_{Y|x})\leq c).$
The derivation indicates that, under reasonable assumptions, an increased number of parameters is advantageous for reducing the $G_{M}(\theta_{1}^{m},q_{Y|x})$.
Suppose $G_{M}(\theta,q_{Y|x})=\mathbb{E}_{X}G(\theta_{k},q_{Y|x})$, meaning $\mathbb{E}_{X}G(\theta_{k},q_{Y|x})\leq\mathbb{E}_{X}G(\theta_{i},q_{Y|x}),\forall i\in[1,m]$, and when the gradient-based algorithm converges to a stationary point with respected to $\theta_{k}$ (i.e., $F(\theta_{k},q_{Y|x})=0$), we obtain:
$\mathrm{fit_{n}}(\ell,f_{\theta})=\sqrt{G_{M}(\theta,q_{Y|x})}.$ (79)
Consequently, we can formulate the following corollary.
Corollary 2 (Effect of parameter quantity).
Under reasonable assumptions, an increase in the number of parameters tends to yield smaller normalized fitting error when using gradient-based algorithms.
6 Bound on expected risk
In this section, we leverage the conclusions of generalization error and fitting error discussed in the previous section to establish an upper bound on the expected risk, offering a new perspective on classification. By applying the triangle inequality of absolute values, we formulate the following theorem.
Theorem 5 (Bound on expected risk).
$R_{\ell}(f_{\theta},\bar{q})$ $\leq\mathrm{gen}(\ell,f_{\theta})+R_{\ell}(f_{\theta},p)$ (80)
$+\sqrt{\mathbb{E}_{X}[F(\theta,q_{Y|x})+G(\theta,q_{Y|x})]\mathbb{E}_{x}[\|\ell(f_{\theta}(x))\|^{2}_{2}]},$
where
$\Pr(\mathrm{gen}(\ell,f_{\theta})\geq\epsilon)\leq\frac{L^{2}C(q)}{2\epsilon^{2}},$ (81)
$L=\|\ell(y,f_{\theta}(x))\|_{\infty},$
$nq\sim\operatorname{Multinomial}(n,\bar{q}).$
Proof.
Utilizing the triangle inequality for absolute value, we obtain
$|R_{\ell}(f_{\theta},\bar{q})-R_{\ell}(f_{\theta},p)|$ $\leq|R_{\ell}(f_{\theta},\bar{q})-R_{\ell}(f_{\theta},q)|$ (82)
$+|R_{\ell}(f_{\theta},q)-R_{\ell}(f_{\theta},p)|$
$=\mathrm{gen}(\ell,f_{\theta})+\mathrm{fit}(\ell,f_{\theta}).$
Eliminating the absolute value signs yields the inequality:
$R_{\ell}(f_{\theta},\bar{q})\leq R_{\ell}(f_{\theta},p)+\mathrm{gen}(\ell,f_{\theta})+\mathrm{fit}(\ell,f_{\theta}).$ (83)
By applying Theorem 4 and 5 to this inequality, we can prove the theorem. ∎
Unlike traditional statistical learning frameworks that use empirical and generalization errors to bound expected error, our theorem employs model risk, generalization error, and fitting error to control the expected error. Contrary to the conventional generalization theories[6, 7] that bound the generalization error by the complexity of the hypothesis space, we utilize the complexity of the data for bounding purpose.
Furthermore, we introduce the term $\mathrm{fit}(\ell,f_{\theta})$ to measure the fitting capacity of hypothesis. This research sheds light on the non-convex optimization mechanism of DNNs by examining the impact of gradient norms and parameter quantities on the fitting error. The primary distinction between the classical bound and our proposed bound on expected risk is illustrated in Figure 1.
Drawing from the findings in Section4 and Section 5, we make the following inferences about the training techniques for DNNs.
Corollary 3.
  1. 1.
    Increasing the number of training samples and decreasing the maximum value of the loss function will help reduce the expected risk.
  2. 2.
    When the complexity $C(q)$ of the training dataset is higher, the weight of the regularization term should also be higher.
  3. 3.
    Lowering the model risk $\mathcal{R}_{M}(f_{\theta})$ is beneficial for reducing expected risk. When cross-entropy is selected as loss function, we have $\mathcal{R}_{M}(f_{\theta})=H(Y|X)$. When keeping $H(Y)$ constant, reducing model risk is equivalent to increasing the mutual information between $X$ and $Y$ where $(X,Y)\sim p$.
  4. 4.
    When using gradient-based algorithms for training, increasing the number of parameters is beneficial for reducing the fitting error and thus the expected risk.
  5. 5.
    A smaller $\lambda_{\max}(H(f_{\theta}(x)))$ is beneficial in reducing the expected risk.
  6. 6.
    Decreasing the empirical risk helps lower the expected risk.
Proof.
Corollary 1 and 2 are the direct consequences of Theorem 5. Since $H(Y|X)=H(Y)-I(X;Y)$, thus when keeping $H(Y)$ constant, decreasing $H(Y|X)$ is equivalent to increasing $I(X;Y)$. Corollary 3 is then proved. Corollary 4 follows directly from Corollary 2. Considering that $\mathbb{E}_{x}[\|\ell(f_{\theta}(x))\|^{2}_{2}]$ in Theorem 5 is equivalent to the empirical risk, then we derive Corollary 5. ∎
7 Empirical Validations
In the preceding section, we theoretically examine the proposed expected risk bound. This section aims to empirically validate its effectiveness. Given the same classification task, the inherent complexity remains constant, thus we concentrate on the fitting error aspect. Within the upper bound of the fitting error, $\mathbb{E}_{X}[G(\theta,q_{Y|x})]$ exhibits stochastic during the gradient descent iterations, while $\mathbb{E}_{X}[F(\theta,q_{Y|x})]$ is consistent with the currently popular flat minima theory under reasonable assumptions, as demonstrated in proposition 4. More importantly, the item $\mathbb{E}_{X}[F(\theta,q_{Y|x})]$ is both controllable and optimizable during the model training process. Consequently, the key to validating the effectiveness of our proposed upper bound lies in confirming the correlation between $\mathbb{E}_{X}[F(\theta,q_{Y|x})]$ and the expected risk.
In this empirical validation, we conducted experiments using 3 distinct neural networks architectures to assess their efficacy in classifying CIFAR-10 dataset, which comprises 50,000 training images and 10,000 testing images averagely distributed in 10 classes. We specifically employed ResNet-18 [58], MobileNetv3 [59] and a custom-designed CNN architecture consisting of five convolutional layers. The architecture of the custom-designed CNN is detailed in Tables 1. We adopted the cross-entropy loss as our optimization objective. All experiments were based on Pytorch. All simulations were conducted on a PC with Intel (R) Core (TM) i7-9700 CPU @ 3.6GHz 32G with two NVIDIA GTX 2080 Ti GPUs. The code for these experiments is available at the provided GitHub repository https://github.com/qibinchuan/Expected-Risk-bound.
Layer Output Shape Implementation Details
Conv1 $8\times 32\times 32$ $3\times 3,8$
Conv2 $16\times 16\times 16$ $1\times 1,16,\text{stride}=2$
Conv3 $16\times 16\times 16$ $3\times 3,16,\text{stride}=2$
Conv4 $32\times 8\times 8$ $3\times 3,32,\text{stride}=2$
Conv5 $64\times 4\times 4$ $3\times 3,64,\text{stride}=2$
Flatten $1024$
Output $10$ Linear
Table 1: Structure of the custom-designed network with 5 conv layers.
In our experiments, we employed the Stochastic Gradient Descent (SGD) optimizer with a weight decay of 0.0005 and a momentum of 0.9. Mini-batch training was conducted with a batch size of 100 for 200 epochs, with an initial learning rate ranging from 0.5 to 0.01. The learning rate was reduced at multiple steps with varying rates. We performed per-pixel mean subtraction, 32 × 32 random cropping after padding with 4 pixels on each side, and horizontal random flipping as data preprocessing and augmentation strategy.
Initially, we trained the models based on the aforementioned configuration. Once the model’s training loss reached stability, indicating minimal change in model’s predictions on the training dataset, we computed the value of $\mathbb{E}_{X}[G(\theta,q_{Y|x})$ and the test accuracy at the end of each epoch, recording a total of 100 epochs. We used the test dataset’s prediction accuracy as a proxy for expected risk, and then calculated the covariance matrices of accuracy and $\mathbb{E}_{X}[G(\theta,q_{Y|x})$ sequences of the three models. It is worth nothing that when training loss stability is stabilized, the model risk is considered constant, and its influence becomes negligible. In addition, due to the significant scale difference between accuracy and $\mathbb{E}_{X}[G(\theta,q_{Y|x})$, we normalized accuracy by dividing it by a constant of 20. This normalization facilitates a clearer visualization of their correlation.
(a) Test Accuracy and $\mathbb{E}_{X}[G(\theta,q_{Y|x})$ sequence of costume designed CNN.
(b) Test Accuracy and $\mathbb{E}_{X}[G(\theta,q_{Y|x})$ sequence of MobileNetv3.
(c) Test Accuracy and $\mathbb{E}_{X}[G(\theta,q_{Y|x})$ sequence of ResNet18.
(d) Heatmap of covariance matrix.
Figure 3: Experimental results for verifying the correlation between expected risk and proposed bound.
The experimental findings, as illustrated in Figure 3, exhibit a strong positive correlation between accuracy and $\mathbb{E}_{X}[G(\theta,q_{Y|x})$ for the same neural network once training loss stabilizes. Conversely, no significant correlation was observed between different networks. This experiment confirms the validity of our proposed upper bound.
8 Conclusion
This study presents a novel information-theoretic analysis of expected risk in supervised classification. To estimate the expected risk, we introduce the concepts of fitting error as well as the model risk. Together with the generalization error, they provide an upper bound on the expected risk. Our analysis reveals that the generalization error can be bounded by the complexity of the data distribution. We also establish a correlation between this complexity and the smoothness of the dataset distribution and suggest that complexity can be used as a measure of the reliability of the training dataset and can guide the setting of hyperparameters for regularization. The derived bound on the fitting error connects the back-propagated gradient, the Neural Tangent Kernel (NTK), and the parameter quantity of model to the fitting error, providing insights into the non-convex optimization mechanism of Deep Neural Networks (DNNs). By leveraging the triangle inequality, we combine complexity and fitting error to derive an upper bound on expected risk. This bound reveals that the expected risk is influenced by both the complexity of the learning task and the fitting errors, and implies the following conclusions under mild conditions.
  • Increasing training samples and decreasing the maximum the loss value contribute to reduction in the expected risk.
  • The expected risk can be effectively controlled by the gradient of the Kullback-Leibler (KL) divergence between the empirical conditional probability distribution and the predictive probability distribution of the model.
  • Gradient-based algorithms can lead to lower expected risk as the number of parameters grows.
  • A smaller maximum eigenvalue of equal-input Neural Tangent Kernel (eNTK), $\lambda_{\max}(H(f_{\theta}(x)))$, which is proved to be equivalent to the flat minima, is advantageous in minimizing the expected risk.
Compared to previous generalization bounds, our bounds are more insightful, as they can help explain many experimental observations. In addition, we have also experimentally verified that there is indeed a significant positive correlation between the proposed upper bounds and the expected risk in real-world scenarios. Thus based on the new upper bound, it becomes possible to achieve better model performance and inspire the development of new algorithms by controlling the key factors in the upper bound.